OpenStudy (anonymous):

which property is illustrated by the statement? 8 + 3.4 = 3.4 + 8 1. associative property of addition. 2. inverse property of multiplication. 3. commutative property of addition. .4. inverse property of addition.

OpenStudy (anonymous):

i think c but this can help you more then i can Commutative Property numbers can be added in any order. symbols=a+b=b+a Associative Property For three or more numbers, their sum is always the same, no matter how the numbers are grouped.symbols=(a+b)+c=a+(b+c) Opposite of a Sum the opposite of the sum of two numbers is equal to the sum of the opposites.symbols=-(a+b)= -a+(b)

jhonyy9 (jhonyy9):

3 is sure